# Modelling Lateral Spread in Wire Flat Rolling

### v1

@@ 2.1 Scaling and non-dimensionalising the equations
The Coulomb friction condition on the roll surface $y=h(x)$ is $\boldsymbol{t\cdot\sigma\cdot n}=\mp\mu\boldsymbol{n\cdot\sigma\cdot n}$, where $\mu=\delta\beta$ is the coefficient of friction, $\boldsymbol{n}=(-\delta\mathrm{d}h/\mathrm{d}x,1,0)/\sqrt{1+\delta^{2}\big{(}\mathrm{d}h/\mathrm{d}x\big{)}^{2}}$ is the unit normal to the roll surface and $\boldsymbol{t}=(1,\delta\mathrm{d}h/\mathrm{d}x,0)/\sqrt{1+\delta^{2}\big{(}\mathrm{d}h/\mathrm{d}x\big{)}^{2}}$ is the unit tangent in the rolling direction. This may be expressed as
$\delta\frac{\mathrm{d}h}{\mathrm{d}x}\big{(}\sigma_{yy}-\sigma_{xx}\big{)}+\!\left(\!1-\delta^{2}\!\left(\frac{\mathrm{d}h}{\mathrm{d}x}\right)^{\!\!2}\right){\sigma}_{xy}\\ =\mp\delta\beta\!\left({\sigma}_{yy}-2\delta\frac{\mathrm{d}h}{\mathrm{d}x}{\sigma}_{xy}+\delta^{2}\!\left(\frac{\mathrm{d}h}{\mathrm{d}x}\right)^{\!\!2}\!{\sigma}_{xx}\right).$ (11)
@@
$\frac{\hat{F}_{\mathrm{A/out}}}{4\hat{\kappa}{\hat{h}^{2}_{\mathrm{A}}}}=\int_{-b_{\mathrm{A/out}}}^{b_{\mathrm{A/out}}}\int_{-h_{\mathrm{A/out}}}^{h_{\mathrm{A/out}}}\sigma_{xx}\,\mathrm{d}y\mathrm{d}z,$ (12)
where $h_{\text{out}}$ and $b_{\text{out}}$ are half of the final thickness and width respectively. Vertical symmetry requires the boundary conditions
$\sigma_{xy}(x,0)$ $=0$ and $v(x,0)$ $=0.$ (13)
The no-flux constraint on the roll surface $y=h(x)$ and wire edge $z=b(x)$ leads to
$v$ $=\delta\frac{\mathrm{d}h}{\mathrm{d}x}u,$ $w$ $=\delta\frac{\mathrm{d}b}{\mathrm{d}x}u.$ (14)
Finally, Integrating the mass conservation law in (8a) from $y=-h(x)$ to $y=h(x)$ and $z=-b(x)$ to $z=b(x)$ yields
$\int_{-b(z)}^{b(z)}\int_{-h(z)}^{h(z)}u\ \mathrm{d}y\mathrm{d}z=4.$ (15)
@@ 2.2 Leading-order asymptotic solution
We now proceed by expanding all variables in an asymptotic series in $\delta$, so that, for example, $\sigma_{xx}=\sigma^{(0)}_{xx}+\delta\sigma^{(1)}_{xx}+\delta^{2}\sigma^{(2)}_{xx}+\cdots$, and collecting like terms. While this formalism could be used to compute the solution to an arbitrary order of accuracy (for example, Erfanian et al. [17] consider up to the second order terms $\sigma^{(2)}_{xx}$ for sheet rolling), here we will only consider the leading order terms. We therefore neglect all terms of $O(\delta)$ or higher to derive the leading-order equations. Substituting the hydrostatic pressure (9) into the yield function (10) at leading order gives
@@
$\sigma^{(0)}_{yy}$ $=-\frac{3p^{(0)}}{2}-\frac{\sqrt{4-3p^{{(0)}^{2}}}}{2},$ (16c)
@@
Using this condition alongside stress solutions (16) implies that $\sigma^{(0)}_{yy}$ and therefore $p^{(0)}$ and $\sigma^{(0)}_{xx}$ are vertically homogeneous which is the same as the slab method’s assumption. We now show that these components are independent of $z$, as well, and only change along the rolling direction, $x$. From (8e) and (8f) $u^{(0)}$ is independent of $y$ and $z$ which along with the continuity equation (8a) and (8g) it can be inferred that $v^{(0)}$ is linear in $y$ and $w^{(0)}$ is linear in $z$. Now the set of equations (8b)–(8d), where their left-hand sides are only functions of $x$ while the stresses are functions of $p^{(0)}$ (equation (16)) shows that in order for a single value for $\lambda^{(0)}$ satisfies them all, $p^{(0)}$ must be a function of $x$ only.
@@
This is the same equation as was derived using the slab method by Kazeminezhad and Taheri [9], although our formulae for $\sigma^{(0)}_{xx}$ and $\sigma^{(0)}_{yy}$ here are different. Equation (19) provides an ordinary differential equation (ODE) in terms of $p^{(0)}$ and $b(x)$, when combined with stress solutions (16).
Another ODE results from the velocity equations. Equations (8e) and (8f) at leading order show that $u^{(0)}$ has no $y$ or $z$ dependence and so is only a function of $x$. The average mass balance (15) then requires that
$u^{(0)}(x)=\frac{1}{b(x)h(x)}.$ (20)
From flow rules (8b) and (8c) and (8d) $\lambda^{(0)}$ is a function of $x$. Therefore, from the tension flow rule in the $y$ direction (8c) along with the symmetry condition (13), the vertical velocity is found as
@@
$\lambda^{(0)}=\frac{\mathrm{d}h/\mathrm{d}x}{bh^{2}s^{(0)}_{yy}},$ (22)
@@
$\lambda^{(0)}=\frac{\mathrm{d}b/\mathrm{d}x}{b^{2}hp^{(0)}}.$ (23)
@@
By solving equations (24) together with (19), $b(x)$ and $p^{(0)}$ are solved. However, in equation (19) the dependency on $b(x)$ can be removed to simplify further; an alternative ODE for $b(x)$ may be found by replacing $u^{(0)}$ from (20) and $\lambda^{(0)}$ from (23) into (8b), as
@@
$4\sigma^{(0)}_{xx}(x=A)$ $=F_{\mathrm{A}}$ and $4bh\sigma^{(0)}_{xx}\Big{|}_{x=1}$ $=F_{\mathrm{out}}.$ (27)
With this, equation (26) can be solved from point A forward with --ve sign and from the exit backwards with $+$ sign. It is interesting to note from equations (24) and (19) that $b(x)$, and consequently the width of contact, $W_{c}$, and the lateral spread of the wire, $W_{t}$, depend on the initial diameter of the wire, reduction ratio, roll radius, and friction coefficient, and not on the roll speed on the material type (e.g. the yield stress $\hat{\kappa}$). The influence of roll speed and material type would become apparent if hardening were incorporated into the analysis, although they likely remain negligible in comparison.
@@ 2.3 Summary and numerical evaluation
The solution for $p^{(0)}$ is chosen to satisfy the forward and backward tension conditions, which are taken to be zero for the results presented below. Therefore, $p^{(0)}$ is solved by integrating equation (26) forward from the entrance with positive sign of friction, and integrating equation (26) backwards from the exit with positive sign of friction, using the Matlab ODE solver ode45 [18]. This is the same solution as the slab method and the two curves thus produced are referred to as the pressure hill and the intersection determines the location of the neutral point.
After solving for stresses, $b(x)$ and consequently the lateral spread, $W_{t}$, is determined by integrating either equation (24) or (25) from point A to the exit. The integration is performed using the Matlab ODE solver ode45, with the initial condition $b(x=A)=1$, which is consistent with the chosen non-dimensionalisation. As detailed in the intoruction, the area of the rectangular cross-section is assumed to be equal to that of the actual wire in the roll gap. As a result, the lateral spread is determined by equating the cross-sectional area of the real shape to that of the rectangle. Therefore
$2\hat{W}_{c}\hat{h}+\pi\hat{h}^{2}$ $=4\hat{b}\hat{h}$$\Rightarrow$ $\hat{W}_{c}$ $=2\hat{b}-\frac{\pi\hat{h}}{2},$ (29)
where $\hat{W}_{c}$ is the width in contact with the roll, and the bulged edges are assumed to form a half-circle. With this, the lateral spread, $\hat{W}_{t}$, is found as
$\hat{W}_{t}=\hat{W}_{c}+2\hat{h}$ (30)
The total computation time varies depending on the tension at the exit, yet in the order of seconds. For zero tension solving stresses and lateral spread take less than a second.
@@ 4.1 Lateral spread
The lateral spread prediction, $\hat{W}_{t}$, for both cases, is plotted in Figure 3 for various diameters and reduction ratios with 50 $\mathrm{mm}$ roll radius. The results are compared with experimental data, Kobayashi’s empirical equation (3), and equation (2) from Kazeminezhad and Taheri [8]. The latter equation is particularly relevant because the wire used in the experiments is stainless steel, which aligns with the materials to which equation (2) applies. In the current model, predictions depend on the friction coefficient, which is challenging to determine experimentally. The experimental data were obtained under non-lubricated test conditions, and a friction coefficient of $\mu=0.25$ has been shown to provide the best agreement with experimental data. Consequently, this value is used for results in Figure 3. The current model fails at larger reductions for $\hat{d}_{0}=2.96\ \mathrm{mm}$ and $\hat{d}_{0}=3.96\ \mathrm{mm}$, as indicated by the unfinished lines, due to a singularity in pressure (see equation 16). Interestingly, these points correspond roughly to the reductions at which experiments encountered difficulties during single-pass rolling and no experimental data is recorded. The model agrees closely with experimental data across all diameters and reduction ratios. Kazeminezhad Equation (2) depends only on the reduction ratio. While it performs well for smaller reduction ratios, it deviates for larger values, with the deviation appearing to depend on wire diameter (e.g., underestimates for $\hat{d}_{0}=5.96\ \mathrm{mm}$ and overestimates for $\hat{d}_{0}=7.96\ \mathrm{mm}$). Kobayashi’s equation depends on roll radius and wire diameter as well as the reduction ratio and seems to be derived for small wire diameter conditions. It underestimates the lateral spread across all wire diameters, particularly at higher reduction ratios, and is the least reliable for predicting the set of experimental data presented in Figure 3.
@@
Figure 3: Lateral spread for stainless steel wires with different diameters, $\hat{d}_{0}$, and reduction ratios, $\hat{d}_{0}/(2\hat{h}_{1})$ from different methods; experimental data, empirical equation 2 from Kazeminezhad and Taheri [8], empirical equation from Kobayashi (3), and the current model. Other parameters used are $(\hat{R},\mu)=(50\,\mathrm{mm},0.25)$.
@@
Some studies suggest that the friction coefficient has a negligible effect on lateral spread [2, 10], attributing this to the movement of lubricant toward the roll edges, driven by the extremely high contact pressure at the entry point. Similarly, the empirical equations (2) and (3) exhibit no dependence on friction. To examine the impact of friction, the ratio of lateral spread to the initial wire diameter is plotted in Figure 4 for two different friction coefficients. In the absence of experimental data under the lubricated condition, the FE results from Vallellano et al. [16] are utilised. The data reported in [16] corresponds to a wire with a 5 $\mathrm{mm}$ diameter, rolled using 75 $\mathrm{mm}$ radius rolls, with no applied forward or backward tension. The yield stress, $\hat{Y}$, is assumed to be constant and equal to 385 $\mathrm{MPa}$ the same as the average value in FE simulations. The Tresca friction law, $\hat{\tau}=m\hat{\kappa}$ with the coefficient of $m=0.25$ is used to generate FE results in [16]. Generally, the relationship $\mu<m/\sqrt{3}$ is used essentially for a simple elastoplastic model [19], where the upper limit, $\mu=0.14$, is used to transfer $m$ into the Coulomb friction coefficient, $\mu$. The data represented by the black line with circles corresponds to the experimental measurements for $\hat{d}_{0}=5.96~{}\mathrm{mm}$ without lubrication ($\mu=0.25$), while the counterpart for $\hat{d}_{0}=5~{}\mathrm{mm}$ is derived from FE simulations using $m=0.25$ (approximated with $\mu=0.14$). To minimise the influence of different wire diameters in this comparison, wires with relatively similar diameters were selected, and the results were scaled by the initial diameter.
@@
Figure 4: Effect of friction coefficient on lateral spread. Results are plotted from different methods; experimental data, FE simulations [16], empirical equation 2 from Kazeminezhad and Taheri [8], empirical equation from Kobayashi (3), and the current model.

### v2

@@ 2.2 Scalings and the non-dimensionalised governing equations
The Coulomb friction boundary condition on the roll surface $y=h(x)$ is $\boldsymbol{t\cdot\sigma\cdot n}=\mp\mu\boldsymbol{n\cdot\sigma\cdot n}$, where $\mu=\delta\beta$ is the coefficient of friction, $\boldsymbol{n}=(-\delta\mathrm{d}h/\mathrm{d}x,1,0)/\sqrt{1+\delta^{2}\big(\mathrm{d}h/\mathrm{d}x\big)^{2}}$ is the unit normal to the roll surface and $\boldsymbol{t}=(1,\delta\mathrm{d}h/\mathrm{d}x,0)/\sqrt{1+\delta^{2}\big(\mathrm{d}h/\mathrm{d}x\big)^{2}}$ is the unit tangent in the rolling direction. This may be expressed as
$\delta\frac{\mathrm{d}h}{\mathrm{d}x}\big(\sigma_{yy}-\sigma_{xx}\big)+\delta\!\left(\!1-\delta^{2}\!\left(\frac{\mathrm{d}h}{\mathrm{d}x}\right)^{\!\!2}\right)\!{\sigma}_{xy}$ (17)
$=\mp\delta\beta\!\left(\!{\sigma}_{yy}-2\delta^{2}\frac{\mathrm{d}h}{\mathrm{d}x}{\sigma}_{xy}+\delta^{2}\!\left(\frac{\mathrm{d}h}{\mathrm{d}x}\right)^{\!\!2}\!{\sigma}_{xx}\!\right)\hfil\text{on }y=h(x).$
@@
$\frac{\hat{F}_{\mathrm{A/out}}}{4\hat{\kappa}\hat{h}_{\!\>\!\mathrm{A}}^{2}}$ $=\int_{-b_{\mathrm{A/out}}}^{b_{\mathrm{A/out}}}\int_{-h_{\mathrm{A/out}}}^{h_{\mathrm{A/out}}}\sigma_{xx}\,\mathrm{d}y\mathrm{d}z,$ atA/exit (18)
where $h_{\text{out}}$ and $b_{\text{out}}$ are respectively half of the final thickness and width, and $h_{A}=1$ and $b_{A}$ are respectively half of the initial thickness and width at location A. Vertical symmetry requires the boundary conditions
$\sigma_{xy}(x,0,z)$ $=0$ and $v(x,0,z)$ $=0.$ (19)
There is no flow of material either through the roll surface $y=h(x)$, nor through the lateral edges of the wire $z=b(x)$. Mathematically, these are imposed as boundary conditions, expressed as
$\boldsymbol{V\cdot n}$ $=0$ $\Rightarrow$ $v$ $=\frac{\mathrm{d}h}{\mathrm{d}x}u$ $\text{on}\,y=h(x),$ (20a)
$\boldsymbol{V\cdot m}$ $=0$ $\Rightarrow$ $w$ $=\frac{\mathrm{d}b}{\mathrm{d}x}u$ $\text{on}\,z=b(x),$ (20b)
where $\boldsymbol{V}=(u,\delta v,\delta w)$ is the velocity made dimensionless with $\hat{U}_{\!\mathrm{A}}$. Finally, integrating the incompressibility equation (14a) vertically from y=h(x)y=-h(x) to $y=h(x)$, laterally from z=b(x)z=-b(x) to $z=b(x)$, and axially from point A to any axial position $x$, using the divergence theorem and applying the boundary conditions (20) yields
$\int_{-b(x)}^{b(x)}\int_{-h(x)}^{h(x)}u\ \mathrm{d}y\mathrm{d}z=4b_{A},$ (21)
where $h=u=1$ and $b=b_{A}$ at point A gives the constant of integration as $4b_{A}$. Equation (21) effectively says that, since no material is lost or gained during rolling and the material is incompressible, the volume rate of flow across the wire cross-section is a constant independent of where along the wire it is measured at.
@@ 2.3 Leading-order asymptotic solution
Following non-dimensionalisation, all terms in the governing equations are now the same size except for the explicit factors of $\delta$, which is small. Following the technique of asymptotics (see, e.g. hinch), we now expand in an asymptotic series in powers of the small parameter $\delta$, such that each successive term is asymptotically smaller than the preceding one. Whilst many complicated expansions are sometimes needed, here a simple expansion in powers of $\delta$ is sufficient, as is the case for sheet rolling (erfanian2025); this is suggested from the fact that the small parameter $\delta$ only appears as integer powers of $\delta$ in the governing equations and boundary conditions above, and may be verified a posteriori once the solution has been calculated. We therefore expand any unknown variable as
$\phi=\phi^{(0)}+\delta\phi^{(1)}+\delta^{2}\phi^{(2)}+O(\delta^{3}),$ (22)
with $\phi^{(0)}$ not identically zero so that $\phi=O(1)$. Here, $\phi$ represents any of the unknown variables $u$, $v$, $w$, $\sigma_{ij}$, $s_{ij}$, $p$ and $\lambda$; for example, $\sigma_{xx}=\sigma^{(0)}_{xx}+\delta\sigma^{(1)}_{xx}+\delta^{2}\sigma^{(2)}_{xx}$$\cdots$. Substituting the asymptotic expansion into the governing equations (13) to (21) allows smaller terms to be neglected relative to dominant ones successively, establishing an ordered hierarchy that leads to an approximate but systematically improvable solution with a quantifiable error. While this formalism could be used to compute the solution to an arbitrary order of accuracy (for example, erfanian2025 consider up to the second order terms $\sigma^{(2)}_{xx}$ for sheet rolling), here we will only consider the leading order terms. We therefore neglect all terms of $O(\delta)$ or higher to derive the leading-order equations. It should be noted that neglecting $O(\delta)$ terms in the expansion does not mean we are setting $\delta$ to zero, but rather that we are calculating the leading-order solution which represents the dominant physical behaviour for small, yet realistic, values of $\delta$, with a relative error of order $O(\delta)$ which could subsequently be corrected for by including higher-order terms. From the hydrostatic pressure (15) at leading order
$\sigma^{(0)}_{yy}=-3p^{(0)}-\sigma^{(0)}_{xx}$ (23)
Substituting (23) into the yield function (16) at leading order gives
@@
$\sigma^{(0)}_{yy}$ $=-\frac{3p^{(0)}}{2}-\frac{\sqrt{4-3p^{{(0)}^{2}}}}{2}.$ (24c)
@@
Using this condition alongside the stress solutions (24) implies that $\sigma^{(0)}_{yy}$ and therefore $p^{(0)}$ and $\sigma^{(0)}_{xx}$ are vertically homogeneous which is the same as the slab method’s assumption. We now show that these components are independent of $z$, as well, and only change along the rolling direction, $x$. From (14e) and (14f), $u^{(0)}$ is independent of $y$ and $z$, which, together with the continuity equation (14a) and (14g), imply that $v^{(0)}$ is linear in $y$ and $w^{(0)}$ is linear in $z$. Now the set of equations (14b)–(14d), where their left-hand sides are only functions of $x$ while the stresses are functions of $p^{(0)}$ (equation (24)) shows that in order for a single value for $\lambda^{(0)}$ satisfies them all, $p^{(0)}$ must be a function of $x$ only.
@@
This is the same equation as was derived using the slab method by kazeminezhad2006pressure, although our formulae for $\sigma^{(0)}_{xx}$ and $\sigma^{(0)}_{yy}$ here are different. Equation (27) provides an ordinary differential equation (ODE) in terms of $p^{(0)}$ and $b(x)$, when combined with stress solutions (24).
Another ODE results from the velocity equations. Equations (14e) and (14f) at leading order show that $u^{(0)}$ has no $y$ or $z$ dependence and so is only a function of $x$. The volume flow rate (21) then requires that
$u^{(0)}(x)=\frac{b_{A}}{b(x)h(x)}.$ (28)
From flow rules (14b)–(14d), $\lambda^{(0)}$ is a function of $x$. Therefore, from the tension flow rule in the $y$ direction (14c) along with the symmetry condition (19), the vertical velocity is found as
@@
$\lambda^{(0)}=\frac{b_{A}\mathrm{d}h/\mathrm{d}x}{bh^{2}s^{(0)}_{yy}}$, (30)
@@
$\lambda^{(0)}=\frac{b_{A}\mathrm{d}b/\mathrm{d}x}{b^{2}hp^{(0)}}.$ (31)
@@
By solving equations (32) together with (27), $b(x)$ and $p^{(0)}$ are solved. However, in equation (27) the dependency on $b(x)$ can be removed to simplify further; an alternative ODE for $b(x)$ may be found by replacing $u^{(0)}$ from (28) and $\lambda^{(0)}$ from (31) into flow rule (14b), as
@@
$4b_{A}\sigma^{(0)}_{xx}(x=\mathrm{A})$ $=F_{\mathrm{A}}$ at point A, (35a)
$4bh\sigma^{(0)}_{xx}(x=1)$ $=F_{\mathrm{out}}$ at exit. (35b)
With this, equation (34) can be solved from point A forward with -ve sign and from the exit backwards with ++ sign. It is interesting to note from equations (32) and (27) that $b(x)$, and consequently the width of contact, $W_{c}$, and the lateral spread of the wire, $W_{t}$, depend only on the initial diameter of the wire, reduction ratio, roll radius, and friction coefficient, and not on the roll speed on the material type (e.g. the yield stress $\hat{\kappa}$). The influence of roll speed and material type would become apparent if hardening were incorporated into the analysis, although they likely remain negligible in comparison. It is worth commenting on the order of magnitude of the error. Based on the scaling in (11) and asymptotic expansion (22),
$\hat{b}=\hat{h}_{\!\>\!\mathrm{A}}b=\hat{h}_{\!\>\!\mathrm{A}}\big(b^{(0)}+\delta b^{(1)}+O(\delta^{2})\big),$ (36)
the solution obtained up to the leading order, $b^{(0)}$, carries a relative error of order $O(\delta)$.
@@ 2.4 Summary and numerical evaluation
The solution for $p^{(0)}$ is chosen to satisfy the forward and backward tension conditions, which are taken to be zero for the results presented below. Therefore, $p^{(0)}$ is solved by integrating equation (34) forward from the entrance with positive sign of friction, and integrating equation (34) backwards from the exit with positive sign of friction, using the Matlab ODE solver ode45 (MATLAB:R2024a_u1). This is the same solution as the slab method, and the two curves thus produced are referred to as the pressure hill, and the intersection determines the location of the neutral point.
After solving for stresses, $b(x)$ and consequently the lateral spread, $W_{t}$, is determined by integrating either equation (32) or (33) from point A to the exit. The integration is performed using the Matlab ODE solver ode45, with the initial condition $b(0)=b_{A}$. For the rolling of a cylindrical wire, we assume $b_{A}=h_{A}=1$, as stated in section 2.1.1 and further justified in A, and consistent with the non-dimensionalisation. Nevertheless, retaining $b_{A}$ explicitly in the governing equations allows the same model to be applied to subsequent roll stands, where the initial cross-section is no longer circular but approximately rectangular with bulged edges. In such cases, location $A$ corresponds to the roll-gap entrance. As detailed in the introduction and shown in figure 2, the area of the rectangular cross-section is assumed to be equal to that of the actual wire in the roll gap. As a result, the lateral spread is determined by equating the cross-sectional area of the real shape to that of the rectangle, with $\hat{W}_{c}$ the width in contact with the roll and $\hat{W}_{t}$ the total lateral spread then given by (5) and (6) respectively.
Recall that $\hat{b}$ in equations (5) and (6) is defined as $\hat{b}=\hat{h}_{\!\>\!\mathrm{A}}b$, where $b$ is determined independently of the material type. As a result, both $W_{c}$ and $W_{t}$ do not depend on the wire material. The total computation time varies depending on the tension at the exit, and is at most in the order of seconds on a standard laptop. For zero tension, solving stresses and lateral spread takes $0.1$ seconds on a single core of a 2019 Intel i7-8665U laptop CPU.
@@ 4.1 Lateral spread
The lateral spread prediction, $\hat{W}_{t}$, for both cases, is plotted in Figure 3 for various diameters and reduction ratios with 50 $\mathrm{mm}$ roll radius.
@@
Figure 3: Lateral spread for stainless steel wires with different diameters, $\hat{d}_{0}$, and reduction ratios, $\hat{d}_{0}/(2\hat{h}_{1})$ from different methods; experimental data, empirical equation 2 from kazeminezhad2005width, empirical equation from Kobayashi (3), and the current model. Other parameters used are $(\hat{R},\mu)=(50\,\mathrm{mm},0.25)$.
The results are compared with experimental data, Kobayashi’s empirical equation (3), and empirical equation (2) due to kazeminezhad2005width. The latter equation is particularly relevant because the parameters in (2) were fitted for stainless steel. In the current model, predictions depend on the friction coefficient, which is challenging to determine experimentally. The experimental data were obtained under non-lubricated test conditions, and a friction coefficient of $\mu=0.25$ has been shown to provide the best agreement with experimental data, consistent with the range typically reported for flat rolling of metal wires (lin2008finite; lee2018development). Consequently, this value is used for the results in Figure 3. The current model fails at larger reductions for $\hat{d}_{0}=2.96\ \mathrm{mm}$ and $\hat{d}_{0}=3.96\ \mathrm{mm}$, as indicated by the unfinished lines, due to a singularity in pressure (see equation 24); interestingly, the experiments also encountered difficulties achieving comparable larger reductions during single-pass rolling in these cases, and so experimental data in these cases is also lacking. Where data is available, the model agrees closely with the experimental data for all diameters and reduction ratios. Kazeminezhad Equation (2) depends only on the reduction ratio. While it performs well for smaller reduction ratios, it deviates for larger values, with the deviation appearing to depend on wire diameter (e.g., underestimates for $\hat{d}_{0}=5.96\ \mathrm{mm}$ and overestimates for $\hat{d}_{0}=7.96\ \mathrm{mm}$). Kobayashi’s equation depends on roll radius and wire diameter as well as the reduction ratio and seems to be derived for small wire diameter conditions. It underestimates the lateral spread across all wire diameters, particularly at higher reduction ratios, and is the least reliable for predicting the set of experimental data presented in Figure 3.
@@
Some studies suggest that the friction coefficient has a negligible effect on lateral spread (carlsson1998contact; kazeminezhad2008error), attributing this to the movement of lubricant toward the roll edges, driven by the extremely high contact pressure at the entry point. Similarly, the empirical equations (2) and (3) exhibit no dependence on friction. To examine the impact of friction, the ratio of lateral spread to the initial wire diameter is plotted in Figure 4 for two different friction coefficients. In the absence of experimental data under the lubricated condition, the FE results from vallellano2008analysis are utilised. The data reported by vallellano2008analysis corresponds to a wire with a 5 $\mathrm{mm}$ diameter, rolled using 75 $\mathrm{mm}$ radius rolls, with no applied forward or backward tension. The yield stress, $\hat{Y}$, is assumed to be constant and equal to 385 $\mathrm{MPa}$ the same as the average value in FE simulations. vallellano2008analysis used the Tresca friction law, $\hat{\tau}=m\hat{\kappa}$ with the coefficient of $m=0.25$ to generate their FE results. Generally, the relationship $\mu<m/\sqrt{3}$ is used essentially for a simple elastoplastic model (zhang2016relationship), where the upper limit, $\mu=0.14$, is used to transfer $m$ into the Coulomb friction coefficient, $\mu$. The data represented by the black line with circles corresponds to the experimental measurements for $\hat{d}_{0}=5.96~\mathrm{mm}$ without lubrication ($\mu=0.25$), while the counterpart for $\hat{d}_{0}=5~\mathrm{mm}$ is derived from FE simulations using $m=0.25$ (approximated by $\mu=0.14$). To minimise the influence of different wire diameters in this comparison, wires with relatively similar diameters were selected, and the results were scaled by the initial diameter.
@@
Figure 4: Effect of friction coefficient on lateral spread. Results are plotted from different methods; experimental data, FE simulations (vallellano2008analysis), empirical equation 2 from kazeminezhad2005width, empirical equation from Kobayashi (3), and the current model.
@@
4.2 Longitudinal spread
The model can be easily used to predict the final length of the wire. Since the volume of metal in the wire remains constant, the wire length varies inversely proportional to the cross-sectional area. The final length of the flattened write, $\hat{L}_{1}$ can therefore be calculated from
$\pi\frac{{\hat{d}_{0}}^{2}}{4}\,\hat{L}_{0}=4\hat{b}_{1}\,\hat{h}_{1}\,\hat{L}_{1},$ (38)
where $\hat{L}_{0}$ is the initial wire length, $\hat{h}_{1}$ is the final half thickness and $\hat{b}_{1}$ the final half width from equation (32).
Figure 5: The initial wire length, along with the final experimental and predicted length for different samples. Each sample is labelled according to its initial wire diameter and the ratio of initial diameter to final thickness (for instance, “2.96–1.25” refers to a wire with an initial diameter of 2.96 mm and a reduction ratio of 1.25). Other parameters used are $(\hat{R},\mu)=(50\,\mathrm{mm},0.25)$.
Figure 5 shows the initial length along with the final experimental and predicted length for different samples. Each sample is labelled according to its initial wire diameter and the ratio of initial diameter to final thickness (for instance, “2.96–1.25” refers to a wire with an initial diameter of 2.96 mm and a reduction ratio of 1.25). Overall, the predicted final lengths show good agreement with the experimental measurements across the tested range of samples: the average absolute error in length prediction is $1.1\,\mathrm{mm}$, and the average relative error in length prediction is $2\,\%$. This confirm that the model reasonably captures the longitudinal deformation associated with lateral spread and reduction during wire rolling.